OpenStudy (anonymous):

What are your solutions when you solve for this System? It's really tricky... x+y=9 y+z=7 x-z=2

OpenStudy (anonymous):

x = 9-y x = 2+z 9-y = 2+z 7 -y = z y+ (7-y) = 7 no solution here need someone to double check

jimthompson5910 (jim_thompson5910):

jayz657, y+ (7-y) = 7 turns into y - y + 7 = 7 which turns into 7 = 7 which is ALWAYS TRUE regardless of what x, y and z are.

jimthompson5910 (jim_thompson5910):

so there are actually an infinite number of solutions

OpenStudy (anonymous):

oh right haha there you go

OpenStudy (anonymous):

So does that mean theres no 3 speciffic answerss?

jimthompson5910 (jim_thompson5910):

depending on how your teacher wants it, the answer is either a) you just say there are infinitely many solutions or b) you say there are infinitely many solutions and that the system is consistent and dependent or c) you write the solution as an ordered triple in the form (x,y,z) where you introduce a free variable (eg: z is a free variable where x and y depend on z)

jimthompson5910 (jim_thompson5910):

If you must answer as described in c) above, then y+z=7 ---> y = 7 - z ------------------------------------------------------- x+y=9 ----> x = 9-y ----> x = 9-(7-z) ----> x = 2 - z So if z is the free variable, then the solution as an ordered triple is ( 2-z, 7-z, z)

jimthompson5910 (jim_thompson5910):

If z = 0, then ( 2-z, 7-z, z) becomes ( 2, 7, 0), which is one solution If z = 1, then ( 2-z, 7-z, z) becomes ( 1, 6, 1), which is another solution If z = 2, then ( 2-z, 7-z, z) becomes ( 0, 5, 2), which is another solution etc etc You can replace z with any real number to get any solution (of the infinitely many solutions)

OpenStudy (anonymous):

ok thnx!

jimthompson5910 (jim_thompson5910):

np